OpenStudy (anonymous):

What are the solutions of the compound inequality? Graph the solutions. –2 < 4x – 10 < 6

OpenStudy (lgbasallote):

\[-2 < 4x - 10 < 6\] add 10 to all sides \[\implies -2 + 10 < 4x - 10 + 10 < 6 + 10\] combine like terms \[\implies 8 <4x < 16\] divide everything by 4 \[\implies \frac 84 < \frac{4x}{4} < \frac{16}{4}\] simplify \[\implies 2 < x < 4\] does that help?

OpenStudy (anonymous):

HMMMM let me seee one moment

OpenStudy (lgbasallote):

sure

OpenStudy (lgbasallote):

tag me if you have follow up questions

OpenStudy (anonymous):

idk what would the answer b ???

OpenStudy (lgbasallote):

??

OpenStudy (lgbasallote):

i just told you...

OpenStudy (anonymous):

o i new that thats what i got !!!